OpenStudy (anonymous):

Points (2, 0) and (0, 3) lie on line k. What is the slope of the line that is perpendicular to k?

OpenStudy (jpsmarinho):

2-0 = m(0-3) m= 2/-3

OpenStudy (jpsmarinho):

y-y0 = m(x-x0)